OpenStudy (rational):

solve Ld^2q/dt^2 + Rdq/dt + q/C = 0

OpenStudy (faiqraees):

are L and R and C variables? are we suppose to find a different equation or find a value for unknown?

OpenStudy (rational):

L R C are constants @FaiqRaees

OpenStudy (rational):

It is a linear second order ordinary differential equation with constant coefficients... I need to find the general solution..

OpenStudy (phi):

don't you use the characteristic equation L r^2 +R r + 1/C =0 to find the two roots r1 and r2 (a bit ugly) and the answer is a linear combination q(t)= c1*exp( r1 t) + c2*exp(r2 t)

OpenStudy (thomas5267):

The solution is a combination of sine and cosine.

OpenStudy (rational):

But my textbook says the solution is a decaying exponential multiplied by cosine @thomas5267

OpenStudy (thomas5267):

Oh crap. Guess \(e^{kx}\) as a solution. Plug it into your equation and solve for k. If the auxillary polynomial has two imaginary roots, then the solution would be the one suggested by your textbook if you rearrange the exponential functions to trig functions.

OpenStudy (rational):

L r^2 +R r + 1/C =0 solving, i get \[r = \dfrac{-R\pm \sqrt{R^2 - 4L/C}}{2L}\]

OpenStudy (thomas5267):

The solution is probably equivalent to the usual solution but it looks weird. In the underdamped case the textbook solution is the usual one but in other cases I don't think so.

OpenStudy (phi):

you can factor out -R/2L if the square root is negative (we need to assume values for the constants) you get an exponential with imaginary component

OpenStudy (thomas5267):

Note that \(\omega_0^2=\dfrac{1}{LC}\).

OpenStudy (thomas5267):

\[ \begin{align*} L\frac{d^2q}{dt^2} + R\frac{dq}{dt} + \frac{q}{C} = 0\\ \frac{d^2q}{dt^2} + \frac{R}{L}\frac{dq}{dt} + \frac{q}{LC} &= 0\\ \frac{d^2q}{dt^2} + \frac{R}{L}\frac{dq}{dt} + \omega_0^2q &= 0\\ \end{align*} \] \[ k^2 + \frac{R}{L}k + \omega_0^2 = 0\\ \begin{align*} k &= \frac{-\frac{R}{L}\pm \sqrt{\frac{r^2}{L^2}-4\omega_0^2}}{2}\\ k &= \frac{-\frac{R}{L}\pm 2\sqrt{\left(\frac{r}{2L}\right)^2-\omega_0^2}}{2}\\ k &= \frac{-\frac{R}{L}\pm 2i\sqrt{\left(\omega_0^2-\frac{r}{2L}\right)^2}}{2}\\ k &= -\frac{R}{2L}\pm i\sqrt{\left(\omega_0^2-\frac{r}{2L}\right)^2} \end{align*}\\ q=c_1e^{-\frac{R}{2L} + i\sqrt{\left(\omega_0^2-\frac{r}{2L}\right)^2}}+c_2e^{-\frac{R}{2L} - i\sqrt{\left(\omega_0^2-\frac{r}{2L}\right)^2}} \] Rearrange to trig form and apply trig identities.

OpenStudy (rational):

Awesome, thanks!

OpenStudy (phi):

I just typed it up, so I may as well post it. the two roots of the characteristic equation are \[r= -\frac{R}{2L} \pm \frac{1}{2L} \sqrt{R^2 -\frac{4L}{C} } \] putting the 1/2L inside the square root, that part becomes \[ \sqrt{\left( \frac{R}{2L}\right)^2 - \frac{1}{LC} }\] replace 1/LC with omega squared factor out a -1 from the square root \[ i \sqrt{ \omega^2-\left(\frac{R}{2L}\right)^2} \] rename the "square root" stuff omega prime. the linear combination of solutions is \[ q = e^{-\frac{R}{2L}t} \left(c_1 e^{i \omega ' t}+ c_2 e^{-i \omega ' t} \right) \] the linear combination of exponentials can be written as a single phase shifted cosine (or sine) (it takes a bit to write out how to do that) \[ q = c e^{-\frac{R}{2L}t} \cos(\omega ' t + \phi) \]

OpenStudy (rational):

I think we're assuming underdamped because \(\omega' \lt \omega\) can be inferred from the textbook... Makes so much sense, thanks a lot!